OpenStudy (anonymous):

Solve for x: x - 0.2x + 0.07(x - 0.2x) = 73.53

OpenStudy (anonymous):

Distributive property and adding like terms ought to clean this up.

OpenStudy (anonymous):

x-.2x+.07x-.014x=73.53 1.486x=73.53 50.8=x

OpenStudy (anonymous):

Soo.. -0.2x + 0.07x + 0.014x = 73.53 then... -0.116?

OpenStudy (anonymous):

@pkjha3105 Your answer must be wrong because it isn't even listed.

OpenStudy (anonymous):

Is \[x \approx85.90\] listed? If so, I will explain how I got it.

OpenStudy (anonymous):

Yes, it is listed

OpenStudy (anonymous):

You actually don't have to use distributive at all; \[x-0.2x+0.07(x-0.2x)=73.53\] In this equation, it helps to write all of the x's that are alone as 1x's. So, \[1x-0.2x+0.07(1x-0.2x)=73.53\] Now solve what's in parentheses first, \[1x-0.2x=0.8x\]So, \[0.8x+0.07(0.8x)=73.53\] Now, \[0.07 \times0.2x=0.056x\]So,\[0.8x+0.056x=73.53\] \[0.856x=73.53\]Divide both sides by 0.856 to get x alone; \[\frac{ 0.856x }{ 0.856 }=\frac{ 73.53 }{ 0.856 }\] \[x = 85.89\] or \[x \approx85.90\] got to love decimals, haha.